Grade 12th passElectrostatics

Two voltmeter have the same range 0-400v . The internal impedence are 32kohm and 18kohm respectively. If they are connected in series and 500v be applied across them . What will be the reading

Pushpam
8 Years agoGrade 12th pass

1 Answer

Eshan
8 Years ago
Dear student,

The 500V input will divide in the ratio of their resistances. Hence the voltage across the two voltmeters and hence their reading would be\dfrac{32}{32+18}400V=256Vand\dfrac{18}{32+18}400V=144V, respectively.

Regards.